OpenStudy (anonymous):

Find the dot product v=3i-5j and w=7i+j

OpenStudy (shamim):

ok. i m trying

OpenStudy (shamim):

\[v.w=(3i-5j).(7i+j)=3 \times 7 (i.i)+3(i.j)-5 \times 7 (j.i)-5(j.j)\]

OpenStudy (shamim):

\[21 \times 1 +3 \times 0 -35 \times 0 -5 \times 1=21-5=16\]

OpenStudy (shamim):

feel free to ask any question

OpenStudy (shamim):

i will b happy to give answer

OpenStudy (anonymous):

thanks:)